# Scalability of Metropolis-within-Gibbs schemes for high-dimensional Bayesian models

### v1

@@ 4.2 Conditionally log-concave distributions
The dependence of the conductance on the dimensionality of the single block is the one we expect from the usual theory on Independent Metropolis-Hastings schemes, i.e. the bound decreases exponentially fast in $d_{i}$. Thus, the best blocking of the variables in terms of conductance is a trade-off between the conditional conductance in (15), which deteriorates by increasing $d_{i}$, and the conductance $\Phi(G)$ of the associated Gibbs sampler, which typically increases by increasing $d_{i}$.
@@
[…]
@@
Also in this case the best blocking scheme, in terms of conductance, is given by a trade-off between the conditional conductance and the behaviour of the Gibbs sampler. However, the dependence on $d_{i}$ is polynomial, rather than exponential as in the case of Proposition 2, which makes RWM more robust to the dimensionality of each coordinate.
@@ 5.1 Approximate conductance and perturbations of Markov operators
with $\mathcal{N}\left(\pi_{1},M\right)$ as in (2). The next theorem shows how to relate $\Delta(P_{1},P_{2},M)$ with the difference between $\Phi_{s}(P_{1})$ and $\Phi_{s^{\prime}}(P_{2})$ for $s^{\prime}$ close to $s$.
@@ 6.1 Gibbs and MwG kernels
Beyond being interesting per se, relating the convergence properties of $P_{J}$ to the ones of $G_{J}$ is theoretically appealing because the latter is potentially much easier to analyse, using for example the dimensionality reduction approach discussed in [3, Lemma 4.2]. On the contrary $P_{J}$ does not enjoy such dimensionality reduction property and thus performing a direct analysis of it is in principle much harder.
@@ 1. $(B1)$
[…]
@@ 3. $(B3)$
Assumptions $(B1)$-$(B3)$ require model (20) to be well-specified, in the sense that the marginal likelihood (26) corresponds to the true data-generating mechanism. Moreover, the global parameters $\psi$ need to be identifiable, as formalized in $(B2)$ and $(B3)$: this guarantees that the posterior distribution on $\psi$ contracts to the true value $\psi^{*}$ at an appropriate rate (thanks to the Bernstein-von Mises Theorem , e.g. [68]). We also consider a set of more technical regularity assumptions on the likelihood $f(y\mid\theta_{j})$, $(B4)$-$(B6)$, which are stated for completeness in Appendix C. Notice that assumptions $(B1)$-$(B6)$ are satisfied by common formulations of model (20), such as the hierarchical normal model or models for binary and categorical data (see e.g. Section $5$ in [3]).
@@ 6.3 Dimension-free mixing of MwG for hierarchical models
The following condition requires the conditional conductance of $P_{J}$ to be bounded away from 00 in a neighbourhood $S^{*}$ of $\psi^{*}$:
@@
The key requirement in (27) is that the kernel $P^{\psi,Y_{j}}$ is well-behaved in a neighbourhood of $\psi^{*}$, uniformly in $Y_{j}\in\mathcal{Y}$. Section 6.5 provides some examples where these conditions are verified.
@@
Consider model (20), kernels $G_{J}$ and $P_{J}$ defined in (22)-(23) and $s\in(0,1/2)$. Then, under assumptions $(B1)$-$(B3)$ and $(C)$, we have
[…]
@@
Consider model (20) and $G_{J}$ defined in (22). Then, under assumptions $(B1)$-$(B6)$, for every $s\in(0,1/2)$ there exists a constant $g(s)>0$ such that
@@
Consider model (20) and $P_{J}$ defined in (23). Under assumptions $(B1)$-$(B6)$ and $(C)$, for every $M\geq 1$ and $\epsilon>0$ there exists $T\left(\psi^{*},\epsilon,M\right)<\infty$ such that
$Q_{\psi^{*}}^{(J)}\left(t_{mix}(P_{J},\epsilon,M)\leq T\left(\psi^{*},\epsilon,M\right)\right)\to 1$ as $\hbox{as }J\to\infty.$
@@
$t_{mix}(P_{J},\epsilon,M)$ $=\mathcal{O}_{P}(1)$ as $\hbox{as }J\to\infty\,,$
@@ 6.5 Models with discrete data
The main requirement in Corollary 2 is Assumption $(C)$, which imposes a bound on the conditional conductance, uniformly over $Y_{j}\in\mathcal{Y}$. A setting where the latter is easily satisfied is given by models with discrete data, where $\mathcal{Y}$ is finite.
@@
[…]
@@
Consider model (30) with likelihood as in (29) and let assumptions $(B1)$-$(B3)$ be satisfied. Let the operator $P_{J}$ be as in (23), where $\kappa\left(P^{\psi,y_{r}}\right)$ is continuous with respect to $\psi$ and $\kappa\left(P^{\psi,y_{r}}\right)>0$ for every $r=0,\dots,m$ and $\psi\in S^{*}$, with $S^{*}$ neighborhood of $\psi^{*}$. Then for every $M\geq 1$ and $\epsilon>0$ there exists $T\left(\psi^{*},\epsilon,M\right)<\infty$ such that
$Q_{\psi^{*}}^{(J)}\left(t_{mix}(P_{J},\epsilon,M)\leq T\left(\psi^{*},\epsilon,M\right)\right)$ $\to 1$ as $\hbox{as }J\to\infty.$
Requiring $\kappa\left(P^{\psi,y_{r}}\right)>0$ for fixed $\psi$ and $y_{r}$ is arguably a mild condition, e.g. it is implied by geometric ergodicity [28].
@@ 6.5.1 Numerical simulations
[…]
@@
[…]
In all the simulations to follows, the covariates used for the data-generation process (excluding the intercept) are independently sampled from a uniform distribution between $-5$ and $5$.
@@ 6.6 Feasible start
[…]
@@
[…]
@@ 7 Bayesian binary regression with unknown prior variance
is strongly log-concave. We denote the condition number of $\pi\left(\bm{\theta}\mid\alpha\right)$ by $c(\alpha)$. Explicit bounds on $c(\alpha)$ are available [18], which however diverge to $\infty$ as $\alpha$ goes to 00. As a result, convergence properties of MCMC algorithms targeting model (36) with fixed $\alpha$ are well understood [14, 17, 18]. However, $\alpha$ is usually unknown in applications and it is typically incorporated as a random variable in the Bayesian model, as in (36). In such cases, the joint posterior $\pi$ is not log-concave on $\mathbb{R}^{d+1}$ and analyzing the convergence of MCMC algorithms performing joint updates of $(\alpha,\bm{\theta})$ can be much harder.
@@
Intuitively, Proposition 6 implies that the following two conditions are sufficient for $P$ to mix fast: the posterior distribution of $\alpha$ concentrates in a region where the condition number $c(\alpha)$ is not too high and the exact GS has a good approximate conductance (i.e. $\Phi_{s}(G)$ is not too close to 00). While providing a lower bound to $\Phi_{s}(G)$ may seem equally challenging as doing it for $\Phi_{s}(P)$, an important simplification (in terms of dimensionality reduction) is available for the exact GS. Let $T=\bm{\theta}^{\top}\Sigma^{-1}\bm{\theta}$ and let $\tilde{\pi}(\text{d}\alpha,\text{d}T):=\mathcal{L}\left(\text{d}\alpha,\text{d}T\mid Y\right)\in\mathcal{P}(\mathbb{R}^{2})$ be the two-dimensional marginal posterior distribution of $(\alpha,T)$, with associated GS kernel $\tilde{G}$. The next lemma provides a lower bound on $\Phi_{s}(G)$ in terms of $\Phi_{s}(\tilde{G})$.
@@
[…]
@@ 8 Data augmentation schemes for discretely-observed diffusions
with $Y_{i}$ denoting the evolution of the diffusion over the interval $(t_{i-1},t_{i})$ and $P_{i}^{\theta}$ a Markov operator on $\mathbb{R}^{(t_{i}-t_{i-1})}$ with invariant distribution $\pi\left(\text{d}Y_{i}\mid\theta,X_{t_{i-1}},X_{t_{i}}\right)$. In words, step (44) requires updating separately the evolution of the diffusion in the $N$ sub-intervals defined by $\mathbb{T}$: this can be done independently thanks to the Markov property of (40). A common choice for $P_{i}^{\theta}$, see e.g. [61], is an Independent Metropolis-Hastings as in Section 4.1, where the new path $Y_{i}^{\prime}$ is sampled according to a Brownian bridge on $[t_{i}-t_{i-1}]$ constrained to be equal to $X_{t_{i-1}}$ and $X_{t_{i}}$ at the endpoints and it is accepted with probability
@@
    There exists $c>0$ such that $X_{t_{i}}\in[-c,c]$ for every $i=1,\dots,N$.
Assumptions $(C1)-(C2)$ are common in this literature (see e.g. [6]) and are satisfied in many interesting cases [50], while $(C3)-(C4)$ allow for technical simplifications: an alternative would be to consider a compact set $K$ which retains most of the posterior mass, as in (9). The next proposition shows that we can then provide a lower bound on the conductance of $P$ in (43) in terms of the corresponding Gibbs sampler.
@@
Let $P$ be as in (43), with $P_{i}^{\theta}$ operators of the Independent Metropolis-Hastings as in (45). Then under assumptions $(C1)-(C4)$ there exists a constant $\kappa=\kappa(c)>0$ such that
@@
Note that, although not explicitly indicated in our notation, the kernels $P$ and $G$ depend on $N$ (as well as on $\mathbb{T}$ and $X_{\mathbb{T}}$). Indeed, the key and non-trivial part in the statement of Theorem 7 is that the constant $\kappa$ in (46) does not depend on $N$, $\mathbb{T}$ and $X_{\mathbb{T}}$. This, for example, implies that, under assumptions $(C1)-(C4)$, the decrease of conductance passing from $G$ to $P$ remains bounded as $N\to\infty$. Extensions of Proposition 7 to the case of approximate conductance, in order to relax assumption $(C3)$, are direct following the approach of the previous sections.
@@ 9 Discussion
The results of Section 3 show that performances of a coordinate-wise scheme can be controlled monitoring two factors: (a) whether conditional updates are close enough to exact sampling (i.e. the conditional conductance is far from zero) and (b) whether the GS itself is a good scheme for the specific sampling problem. For example, the first inequality in (9) bounds the slow-down of a general coordinate-wise sampler relative to GS (measured in terms of reduction of conductance) with a multiplicative term of the form $\min_{i\in\{1,\dots,d\}}\,\kappa_{i}(P_{i},\mathcal{X})$, defined in (7). An interesting aspect of such a ‘slow-down’ factor is that it depends on $d$ only through the minimum operation. Thus, if the GS updates are replaced by moderately good MH conditional updates (e.g. ones with conductance uniformly bounded away from 00) the resulting MwG sampler will incur in a slow-down that is constant with respect to $d$. This observation agrees with the observation that, provided the full-conditional distributions are well-behaved, MwG tends to perform similarly to the corresponding GS (see e.g. Figure 1), even when $d$ is large.
Note that the results of Section 3, in their current form, strongly relies on the random-scan architecture, where at each iteration a randomly chosen block is updated. Extensions to other popular orderings, e.g. deterministic-scan ones, are less trivial than one might expect. In particular, naive applications of the techniques employed in this paper would lead to conductance bounds depending on multiplicative factors of the form $\prod_{i=1}^{d}\kappa_{i}(P_{i},K)$, which scale exponentially badly with $d$. Developing tight bounds for deterministic-scan coordinate-wise samplers is an interesting direction for future work.

### v2

@@ 4.2 Conditionally log-concave distributions
The dependence of the conductance on the dimensionality of the single block is the one we expect from the usual theory on Independent Metropolis-Hastings schemes, i.e. the bound decreases exponentially fast in $d_{i}$. Thus, the best blocking of the variables in terms of conductance is a trade-off between the conditional conductance in (15), which deteriorates by increasing $d_{i}$, and the conductance $\Phi(G)$ of the associated Gibbs sampler, which typically increases by increasing $d_{i}$ (see e.g. [43, Thm.1] or [66, Sec.2.4]).
@@
$Q_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})=N\left(y_{i};x_{i},\sigma_{i}^{2}(\textbf{x}_{-i})\mathbb{I}_{d_{i}}\right)\text{d}y_{i},\quad\sigma_{i}^{2}(\textbf{x}_{-i})=\frac{\eta}{d_{i}L_{i}(\textbf{x}_{-i})},$ (16)
@@
Also in this case the best blocking scheme, in terms of conductance, is given by a trade-off between the conditional conductance and the behaviour of the Gibbs sampler. However, the dependence on $d_{i}$ is polynomial, rather than exponential as in the case of Proposition 2, which makes RWM more robust to the dimensionality of each coordinate. Positive semi-definiteness of MwG operators as in (16) follows again by Lemma $3.1$ in [5].
@@ 5.1 Approximate conductance and perturbations of Markov operators
with $\mathcal{N}\left(\pi_{1},M\right)$ as in (2). Note that the above discrepancy is not symmetric, i.e. $\Delta(P_{1},P_{2},M)\neq\Delta(P_{2},P_{1},M)$. The next theorem shows how to relate $\Delta(P_{1},P_{2},M)$ with the difference between $\Phi_{s}(P_{1})$ and $\Phi_{s^{\prime}}(P_{2})$ for $s^{\prime}$ close to $s$.
@@ 6.1 Gibbs and MwG kernels
Beyond being interesting per se, relating the convergence properties of $P_{J}$ to the ones of $G_{J}$ is theoretically appealing because the latter is potentially much easier to analyse, using for example the dimensionality reduction approach discussed in [4, Lemma 4.2]. On the contrary $P_{J}$ does not enjoy such dimensionality reduction property and thus performing a direct analysis of it is in principle much harder.
@@ 1. $(B1)$
    There exists $\psi^{*}\in\mathbb{R}^{D}$ such that $Y_{j}\overset{\text{iid}}{\sim}Q_{\psi^{*}}$ for $j=1,2,\dots$. Moreover the map $\psi\to g(\cdot\mid\psi)$ is one-to-one and the map $\psi\to\sqrt{g(x\mid\psi)}$ is continuously differentiable for every $x$. Finally, the prior density $p_{0}$ is continuous and strictly positive in a neighborhood of $\psi^{*}$.
@@ 3. $(B3)$
Assumptions (B1)(B1)-$(B3)$ require model (20) to be well-specified, in the sense that the marginal likelihood (26) corresponds to the true data-generating mechanism. Moreover, the global parameters $\psi$ need to be identifiable, as formalized in $(B2)$ and $(B3)$: this guarantees that the posterior distribution on $\psi$ contracts to the true value $\psi^{*}$ at an appropriate rate (thanks to the Bernstein-von Mises Theorem , e.g. [78]). We also consider a set of more technical regularity assumptions on the likelihood $f(y\mid\theta_{j})$, (B4)(B4)-$(B6)$, which are stated for completeness in Appendix E. Notice that assumptions (B1)(B1)-$(B6)$ are satisfied by common formulations of model (20), such as the hierarchical normal model or models for binary and categorical data (see e.g. Section $5$ in [4]).
@@ 6.3 Dimension-free mixing of MwG for hierarchical models
The following condition requires the conditional conductance of $P_{J}$ to be bounded away from 0 in a neighbourhood $S^{*}$ of $\psi^{*}$:
@@
The key requirement in (27) is that the kernel $P^{\psi,Y_{j}}$ is well-behaved in a neighbourhood of $\psi^{*}$, uniformly in $Y_{j}\in\mathcal{Y}$. Indeed, as mentioned after (25), the update on the (low dimensional vector) $\psi$ can be often performed in closed form, so that the second requirement of (27) is automatically satisfied. Section 6.5 provides some examples where these conditions are verified.
@@
Consider model (20), kernels $G_{J}$ and $P_{J}$ defined in (22)-(23) and $s\in(0,1/2)$. Then, under assumptions (B1)(B1)-$(B3)$ and $(C)$, we have
$Q_{\psi^{*}}^{(J)}\left(\Phi_{s}(P_{J})\geq\frac{\kappa^{2}}{8}\Phi_{s}(G_{J})\right)\to 1$ $\hbox{ as }J\to\infty\,.$
@@
Consider model (20) and $G_{J}$ defined in (22). Then, under assumptions (B1)(B1)-$(B6)$, for every $s\in(0,1/2)$ there exists a constant $g(s)>0$ such that
@@
Consider model (20) and $P_{J}$ defined in (23). Under assumptions (B1)(B1)-$(B6)$ and $(C)$, for every $M\geq 1$ and $\epsilon>0$ there exists $T\left(\psi^{*},\epsilon,M\right)<\infty$ such that
$Q_{\psi^{*}}^{(J)}\left(t_{mix}(P_{J},\epsilon,M)\leq T\left(\psi^{*},\epsilon,M\right)\right)\to 1$ $\hbox{as }J\to\infty.$
@@
$t_{mix}(P_{J},\epsilon,M)$ $=\mathcal{O}_{P}(1)$ $\hbox{as }J\to\infty\,$,
@@ 6.5 Models with discrete data
The main requirement in Corollary 2 is Assumption $(C)$, which imposes a bound on the conditional conductance, uniformly over $Y_{j}\in\mathcal{Y}$. A setting where the latter is easily satisfied is given by models with discrete data, where $\mathcal{Y}$ is finite, provided exact sampling for $\psi$ is available (as will be the case in all the numerical examples of this paper).
@@
$Y_{j}\mid\theta_{j}\sim f(y\mid\theta_{j})\,,\quad\theta_{1},\dots,\theta_{J}\mid\mu,\tau\overset{\text{iid}}{\sim}N(\mu,\tau^{-1})\,,\quad(\mu,\tau)\sim p_{0}(\cdot)\,.$ (30)
@@
Consider model (30) with likelihood as in (29) and let assumptions (B1)(B1)-$(B3)$ be satisfied. Let the operator $P_{J}$ be as in (23), where $\kappa\left(P^{\psi,y_{r}}\right)$ is continuous with respect to $\psi$ and $\kappa\left(P^{\psi,y_{r}}\right)>0$ for every $r=0,\dots,m$ and $\psi\in S^{*}$, with $S^{*}$ neighborhood of $\psi^{*}$. Then for every $M\geq 1$ and $\epsilon>0$ there exists $T\left(\psi^{*},\epsilon,M\right)<\infty$ such that
$Q_{\psi^{*}}^{(J)}\left(t_{mix}(P_{J},\epsilon,M)\leq T\left(\psi^{*},\epsilon,M\right)\right)$ $\to 1$ $\hbox{as }J\to\infty.$
Requiring $\kappa\left(P^{\psi,y_{r}}\right)>0$ for fixed $\psi$ and $y_{r}$ is arguably a mild condition, e.g. it is implied by geometric ergodicity [70, 33].
@@ 6.5.1 Numerical simulations
$Y_{ij}\mid\bm{\theta}_{j},X_{i}\overset{\text{ind.}}{\sim}\,\text{Bernoulli}\left(\frac{e^{X_{i}^{\top}\bm{\theta}_{j}}}{1+e^{X_{i}^{\top}\bm{\theta}_{j}}}\right),\quad\bm{\theta}_{j}\mid\bm{\mu},\bm{\tau}\overset{\text{iid}}{\sim}\otimes_{k=1}^{\ell}N(\mu_{k},\tau_{k}^{-1})\,,\quad(\bm{\mu},\bm{\tau})\sim p_{0}(\cdot),$ (31)
@@
$\mu_{k}\mid\tau_{k}\overset{\text{ind.}}{\sim}N(0,10^{3}/\tau_{k})\,,\quad\tau_{k}\overset{\text{iid}}{\sim}\text{Gamma}(1,1),\quad k=1,\dots,\ell.$ (32)
In all the simulations to follows, the covariates used for the data-generation process (excluding the intercept) are independently sampled from a uniform distribution between 5-5 and $5$.
@@ 6.6 Feasible start
$Q_{\psi^{*}}^{(J)}\left(t_{mix}(P_{J},\epsilon,\mu_{J})\leq T\left(\psi^{*},\epsilon\right)\right)\to 1$ $\hbox{ as }J\to\infty\,.$ (34)
@@
Notice that $\mu_{J}$ defined in (33) is not in general a warm start, i.e. we cannot guarantee $\mu_{J}\in\mathcal{N}\left(\pi_{J}\left(\text{d}\psi,\text{d}\bm{\theta}\right),M\right)$ for a fixed $M$ that does not grow with $J$. However in the proof of Proposition 5 it is shown that $\mu_{J}$ is “close enough” to a warm distribution in total variation distance, which suffices for our purposes.
@@ 7 Bayesian binary regression with unknown prior variance
is strongly log-concave. We denote the condition number of $\pi\left(\bm{\theta}\mid\alpha\right)$ by $c(\alpha)$. Explicit bounds on $c(\alpha)$ are available [21], which however diverge to $\infty$ as $\alpha$ goes to 0. As a result, convergence properties of MCMC algorithms targeting model (36) with fixed $\alpha$ are well understood [16, 20, 21]. However, $\alpha$ is usually unknown in applications and it is typically incorporated as a random variable in the Bayesian model, as in (36). In such cases, the joint posterior $\pi$ is not log-concave on $\mathbb{R}^{d+1}$ and analyzing the convergence of MCMC algorithms performing joint updates of $(\alpha,\bm{\theta})$ can be much harder.
@@
[…]
@@
The main intuition underlying Lemma 4 is that the update of $\alpha$ given $\bm{\theta}$ depends only on $T$, i.e. $\pi\left(\text{d}\alpha\mid\bm{\theta}\right)=\pi\left(\text{d}\alpha\mid T\right)$ for $T=\bm{\theta}^{\top}\Sigma^{-1}\bm{\theta}$, which implies that $G$ and $\tilde{G}$ have closely related convergence properties. See Lemma 9 and the proof of Lemma 4 in the supplement for more details and [65, 18, 58, 4] for results similar in spirit.
@@ 8 Data augmentation schemes for discretely-observed diffusions
with $Y_{i}$ denoting the evolution of the diffusion over the interval $(t_{i-1},t_{i})$ and $P_{i}^{\theta}$ a Markov operator on $\mathbb{R}^{(t_{i}-t_{i-1})}$ with invariant distribution $\pi\left(\text{d}Y_{i}\mid\theta,X_{t_{i-1}},X_{t_{i}}\right)$. In words, step (44) requires updating separately the evolution of the diffusion in the $N$ sub-intervals defined by $\mathbb{T}$: this can be done independently thanks to the Markov property of (40). A common choice for $P_{i}^{\theta}$, see e.g. [69], is an Independent Metropolis-Hastings as in Section 4.1, where the new path $Y_{i}^{\prime}$ is sampled according to a Brownian bridge on $t_{i}-t_{i-1}]$ constrained to be equal to $X_{t_{i-1}}$ and $X_{t_{i}}$ at the endpoints and it is accepted with probability
@@
    There exists $c>0$, not depending on $N$, such that $X_{t_{i}}\in[-c,c]$ for every $i=1,\dots,N$.
Assumptions (C1)(C2)(C1)-(C2) are common in this literature (see e.g. [9]) and are satisfied in many interesting cases [57]. On the other hand, assumptions $(C3)$ and especially $(C4)$ are arguably strong assumptions, which allow for technical simplifications but may not be satisfied in many practical scenarios. An alternative would be to consider a compact set $K$ which retains most of the posterior mass, as in (52). The next proposition shows that we can then provide a lower bound on the conductance of $P$ in (43) in terms of the corresponding Gibbs sampler.
@@
Let $P$ be as in (43), with $P_{i}^{\theta}$ operators of the Independent Metropolis-Hastings as in (45). Then under assumptions (C1)(C4)(C1)-(C4) there exists a constant $\kappa=\kappa(c)>0$ such that
@@
Note that, although not explicitly indicated in our notation, the kernels $P$ and $G$ depend on $N$ (as well as on $\mathbb{T}$ and $X_{\mathbb{T}}$). Indeed, the key and non-trivial part in the statement of Theorem 7 is that the constant $\kappa$ in (46) does not depend on $N$, $\mathbb{T}$ and $X_{\mathbb{T}}$. This, for example, implies that, under assumptions (C1)(C4)(C1)-(C4), the decrease of conductance passing from $G$ to $P$ remains bounded as $N\to\infty$. Extensions of Proposition 7 to the case of approximate conductance, in order to relax assumption $(C3)$, are direct following the approach of the previous sections.
@@ 9 Discussion
The results of Section 3 show that performances of a coordinate-wise scheme can be controlled monitoring two factors: (a) whether conditional updates are close enough to exact sampling (i.e. the conditional conductance is far from zero) and (b) whether the GS itself is a good scheme for the specific sampling problem. For example, the first inequality in (52) bounds the slow-down of a general coordinate-wise sampler relative to GS (measured in terms of reduction of conductance) with a multiplicative term of the form $\min_{i\in\{1,\dots,d\}}\,\kappa_{i}(P_{i},\mathcal{X})$, defined in (7). An interesting aspect of such a ‘slow-down’ factor is that it depends on $d$ only through the minimum operation. Thus, if the GS updates are replaced by moderately good MH conditional updates (e.g. ones with conductance uniformly bounded away from 0) the resulting MwG sampler will incur in a slow-down that is constant with respect to $d$. This observation agrees with the observation that, provided the full-conditional distributions are well-behaved, MwG tends to perform similarly to the corresponding GS (see e.g. Figure 1), even when $d$ is large.
Note that the results of Section 3, in their current form, strongly relies on the random scan architecture, where at each iteration a randomly chosen block is updated. Extensions to other popular orderings, e.g. deterministic-scan ones, are less trivial than one might expect. In particular, naive applications of the techniques employed in this paper would lead to conductance bounds depending on multiplicative factors of the form $\prod_{i=1}^{d}\kappa_{i}(P_{i},K)$, which scale exponentially badly with $d$. Developing tight bounds for deterministic-scan coordinate-wise samplers is an interesting direction for future work.